OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3

OpenStudy (anonymous):

The slope between two points is\[slope = \frac{y_2-y_1}{x_2-x_1}\] Where x1=-1, y1=-5, x2=7, y2=-3

OpenStudy (anonymous):

m=(-5+3)/(-1-7)=-2/-8=1/4